OpenStudy (anonymous):

The freezing point depression of a .100m MgSO4 solution is 0.225C. Determine the experimental van't hoff factor of MgSO4 at this concentration. Please help! I know the answer but have no idea how to do it.

OpenStudy (aaronq):

do you know the equation to use?

OpenStudy (anonymous):

deltaT=i(Kf)(m)??

OpenStudy (aaronq):

yeah, just plug in and solve

OpenStudy (anonymous):

ok, so would it be: 0.225=2(1.86)m?

OpenStudy (aaronq):

nope you're solving for i

OpenStudy (anonymous):

ok, so....0.225=i(1.86)(.100)??

OpenStudy (aaronq):

yes

OpenStudy (anonymous):

ooohhh ok.